OpenStudy (anonymous):

find an equation of the circle that has center (-2,5) and passes through (2,-2)

OpenStudy (mathmate):

If (x0,y0) is the centre of the circle with radius r, then the equation would be (x-x0)^2+(y-y0)^2=r^2 Here you know (x0,y0) but need r. So substitute (x,y)=(2,-2), and (x0,y0)=(-2,5) in the above equation of the circle to find r^2 to complete the answer.

OpenStudy (anonymous):

so 65?

OpenStudy (mathmate):

Yes, 65 equals r^2. So can you now show the complete equation of the circle?

OpenStudy (anonymous):

r^2=65?

OpenStudy (mathmate):

Use " the equation would be (x-x0)^2+(y-y0)^2=r^2" and substitute x0, y0 and r^2 using known values.

OpenStudy (anonymous):

how would that look/

OpenStudy (anonymous):

this is the part I have trouble with

OpenStudy (mathmate):

The general equation of a circle with centre (x0,y0) and radius r is given by: (x-x0)^2+(y-y0)^2=r^2 Since you are given the coordinates of the centre (x0,y0), and you have calculated r^2, you answer to the question will be an equation in x and y, but using numeric values for x0, y0 and r^2.

OpenStudy (anonymous):

so its 2(-2)^2+-2(5)^2=65^2

OpenStudy (anonymous):

2(-2(0)^2+-2(0)^2=65

OpenStudy (anonymous):

like that?

OpenStudy (mathmate):

Since we don't know x and y, they stay in the equation. We only substitute x0, y0 and r^2. Can you give it another try?

OpenStudy (anonymous):

x-(-2)^2+y-(5)^2=65

OpenStudy (mathmate):

It is almost good, except that the parentheses are not perfect. It should read: (x-(-2))^2+(y-5)^2=65 which simplifies to (x+2)^2+(y-5)^2=65

OpenStudy (anonymous):

thank you!

OpenStudy (mathmate):

yw! :)